David Is 6 years older than his sister Delia. In 10 years the sum of There age 52. how old are David and Delia
today

Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52788)

.
If David is x years old, then Delia is (x-6) years old.


In 10 years David will be  (x+10)  years old; Delia will be ((x-6)+10) = (x+4) years old.


The condition says


    (x+10) + (x+4) = 52.


It implies


    2x + 14 = 52,

    2x      = 52 - 14 = 38

     x      = 38/2 = 19.


ANSWER.  David's age is 19 years.  Delia is 19-6 = 13 years old.

Solved.

----------------

There is a bunch of lessons on age word problems in this site:
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problems for mental solution
    - One unusual age word problem
    - Age problem for the day of April, 1
    - OVERVIEW of lessons on age problems

Read them attentively and become an expert in this field.


Answer by greenestamps(13200)



Something like this, if you want/need to use formal algebra....

L = Delia now
L+6 = David now

L+10 + Delia 10 years from now
L+16 = David 10 years from now

(L+10)+(L+16) = 52

Solve using basic algebra.

Or informally....

Sum of their ages 10 years from now is 52.
So sum of their ages now is 52-20 = 32.
Subtract 6 years from David's age to make them the same age; the sum of their ages is now 32-6 = 26.
Then Delia's age is 26/2 = 13.
Add the 6 years back on to David's age to find out he is 13+6 = 19.